OpenStudy (anonymous):

-2 is a root of 3x + 2 = x - 2

OpenStudy (anonymous):

3x+2=x-2 = 3x-x+2+2=0 = 2x+4 = 0 now putting the value of -2 in equation u get : 0 = 0 hence -2 is root of equation